17. If a stock declines from $40 per share to $35 in one year, what is the HPR? (5)
Answer #1

17) HPR ( Holding Period Return)

HPR is total return from Income and assets appreciation over a period of time expressed as a percentage

HPR = Income +(end price value - initial price value)/initial price value × 100

Income - 0

End price value of stock is $35

Initial price value of stock is $40

HPR = 0+(35-40)/40 × 100

HPR= 0+(-5)/40 ×100 = -5/40 ×100= - 12.5%

HPR is -12.5%